Statistics from a county public health nurse is provided below. Assist her in completing her report by determining the rate for each vital statistic based on the following information.

if a rate can not be calculated, note what information is missing. Provide the rate or information needed in column J, coinciding with the information on that rate. For this county, the midyear population is 500,000, the estimated female population at that time was 75,000.

*Note that all rates are multiplied by 1,000 except Maternal Mortality is by a factor of 10,000.

Available information for the year

1. Births: 750
Answer #1

Total mid year population is 500000

Number of female population is 75000

Number of births - 750

The missing data is about number of maternal deaths ,if it was available then we can find the maternal mortality rate using number of births .

We can find the crude rate with the above data

Crude rate = number of births / total population × 1000

= 750/ 500000 × 1000

= 75/50

= 3/2 = 1.5 births for every 1000 people